NAOJ GW Elog Logbook 3.2

gain 2, 20 dB attenuation, 30 Hz lowpass

[Aritomi, Eleonora, Matteo]
First we checked alignment of LO inside AMC. The alignment was very very bad and we couldn't find any resonance. So we decided to remove a front mirror of AMC and aligned from the scratch. While we were doing alignment, we found that a mirror just in front of AMC was loose. That's why we sometimes lost alignment of AMC suddenly. After we fixed it and aligned AMC, we recovered the alignment of LO inside AMC.
For LO, peak is 8.16V and mismatching is 5.6mV+4.8mV+3.2mV. The mode matching is 99.8%.
For BAB, we moved a lens in s pol OPO trans path to improve mode matching. The lens position was 89.5 mm before and now it's 99 mm. Peak is 232mV and mismatch is 9.2mV. The mode matching is 96.2%.
We'll check if the alignment keeps fine or not tomorrow.

[Miyakawa-san, Matteo, Eleonora]
Yesterday, in the afternoon, Miyakawa-san came to NAOJ to finalize the installation of KAGRA DGS.
1) He set up the router to make possible to access the computer from outside. Actually, we couldn't select a fix IP address for the moment so in case the assigned (DHCP) address changes we shoud use the new one for the remote access. The current one is 133.40.117.62.
2) We powered the DAC, AI and AA modules with +/-18 V. The total current needed for the 3 moduels is about 1.2 A so we had to use two power supplies.
4) We connected the AI and AA modules to DAC and ADC with Dsub9 cables I borrow in Kamioka last week.
3) We created a test simulink model and started to do basic tests on ADC and DAC channes. No major problems have been found so far.
Tests will continue in the next days.

[Aritomi, Miyakawa, Oshino (remotely)]
Currently IP address of a computer and a router is obtained by DHCP, so first we have to fix an IP address. To change the network setting of a computer, we edited /etc/network/interfaces, but we can't fix an IP address so far. Miyakawa-san will take over this Sunday.

Participant: Aritomi, Matteo, and Yuhang
We put two ND filters between the last steering mirror before OPO and the last lens before OPO. They are OD0.1 and OD 0.4. So it gives 10^-0.5 = 0.316 factor to both p-pol and coherent control beam.
The reason is to reduce the noise of coherent control OPO transmission coupled into the homodyne detector.

Participant: Marco and Yuhang
After lock coherent control PLL with a 7MHz frequency offset with the main laser, we tried to demodulate the OPO reflection signal with 14MHz. And the OPO transmission is going to the homodyne detector. Then there is demodulation of 7MHz(7MHz local oscillator was connected).
After the demodulation of each signal, we make them go through the filter separately. And then goes to the green phase shifter and the infrared phase shifter separately. After close these two loops, we can basically lock them.
We scanned the green and infrared phase. The error signal for the green phase is around 70mV p-p while 700mV p-p for IR(green power now is 50mW). This may be the reason why we can lock the IR phase much better.
I measured the optical-mechanical transfer function of green phase part. It shows wired behavior. However, we checked together and the measurement strategy looks reasonable. The result is shown in the attached picture and should be further investigated.
Locking condition, for now | |
green phase | Low pass 3Hz Gain 100 |
IR phase | Low pass 1Hz Gain 50 |

Participant: Marco and Yuhang
Since the locking of coherent control PLL plays a very important role in getting a coherent control error signal. When we want to perform coherent control, we suffered a lot from the bad locking condition.
Then we start to consider why we cannot lock coherent control PLL. At some point, we realize, even we used an attenuator to reduce minor peaks. But the PLL system still tries to bring it to these minor peaks. So we decide to remove the attenuator. Now the situation is that the beat note comes from fiber PD and then has an amplification of 18dB.
In this case, we can lock coherent control PLL much better. The locking scheme(named as PLL_CC20190221) is saved in the default folder of all the PLL settings. And the parameters are listed in the attached figure.

From the datasheet of ADF4002 (page 3), there is REFIN input frequency limit from 20MHz to 300MHz. However, we used to send a 7MHz reference signal into REFIN. So in this case, it seems the old locking scheme of coherent control PLL can be improved.
So we decide to use 21MHz REFIN and divide it by 3. This 3 is the value of R.
Today we tried this new locking scheme. However, it still didn't work.
What we observed was PLL locked on the beat note while it goes away easily(Can we put an integrator?). By turning on slow, it can go back to locking point. But there is always overshot(still high gain?).
However, sometimes, we can lock it successfully. So it seems the shape of the locking filters should be improved. Now it works like a not optimized control loop. Maybe we should measure the open loop transfer function of it?

Since we want a reasonably large coherent control error signal. We exchange 'Demod CC' amplification channel with 'EOM SHG+IRMC' channel.
Reason: SHG EOM channel has an amplification factor of 20dB but we are using a 12dB attenuator to reduce sideband amplitude. This means 8dB of amplification is enough.
While CC DEMOD channel had amplification of 14dB. Actually, the more the better. So we decide to exchange it with SHG EOM channel.
After the exchange, we put 6dB attenuator for SHG EOM.
In practice, we could use the AOM channel which has a 37dB amplification. However, I tested it and the AOM channel seems broken.
In the end, the coherent control error signal becomes around 60mV peak to peak.


In order to have the HeNe with the same size as the pump, we want to add 2 lenses between the last 2 mirrors of the HeNe path before the sample.
Attach a picture with the distances from the sample and the screenshot of the Jammat simulation to find the focal lengths of the 2 lenses.

using the calibration reported in entry 1221 we did 2 rectangular perpendicular maps (xz and yz) of the sample S6


Current (A) | Temperature (deg) | |
CC | 1.183 | 38.16 |
P pol | 1.338 | 32.5 |

After the sample N1 was mounted, a set of measurements was performed, with a pump power of 10.6 W on the sample.
We found that a parasitic reflection from the red probe is propagating back towards the IR line. We fixed the problem by slightly moving the diaphragm in front of the chopper. This does not affect the power measured by the power meter
Calibration with reference sample.
Position of the detection stage: 70 mm
Pump power: 34 mW * sqrt(0.55) = 0.02521 W
Probe Beam alignment: DC probe maximized at 4.85 V
AC signal at scan center = 0.08 V
Z stage alignment: surface peak maximum at Z =38.85
R = AC/DC/Abs/P = 0.63
Sample N1
DC level 4.92 V
P = 10 W (current = 7.5 A)
Power transmitted 9 W
Power incident 10.54W
T=9/10.54=0.854
Previous results confirmed (Fig 1). All OK.
Sample S1 mounted.
Transmitted power = 9 W
Preliminary absorption estimate:
AC/DC/P/sqrt(0.85)/R*3.34= 0.00175/4.25/10/sqrt(0.85)/0.63*3.34 = 237 ppm/cm
but DC was not maximized... to be done again
Figure 2
Sample S2 Mounted
Transmitted power = 9 W
Preliminary absorption estimate:
AC/DC/P/sqrt(0.85)/R*3.34= 0.001/4.95/10/sqrt(0.85)/0.63*3.34 = 116 ppm/cm
Figure 3
Sample S3 Mounted
Transmitted power = 9 W
Preliminary absorption estimate:
AC/DC/P/sqrt(0.85)/R*3.34= 47 - 163 ppm/cm
Figure 4
Sample S4 Mounted
Transmitted power = 9 W
Preliminary absorption estimate:
AC/DC/P/sqrt(0.85)/R*3.34= 81 - 406 ppm/cm
Figure 5
Sample S5 Mounted
Transmitted power = 9 W
Preliminary absorption estimate:
AC/DC/P/sqrt(0.85)/R*3.34= 115 - 287 ppm/cm
Figure 6
Sample S6 Mounted
Transmitted power = 9 W
Preliminary absorption estimate:
AC/DC/P/sqrt(0.85)/R*3.34= 80 ppm/cm
Figure 7



Participant: Marco and Yuhang
Since we suspected the OPO reflection signal may be too much and saturate PD. We decided to use TAMA PD. Because:
1. It separates DC and AC. So there is less probability of saturation.
2. TAMA PD amplifies signal around 15MHz. We can see that it also amplifies 14MHz quite a lot from the attached figure.
However, even after replacement, we cannot see useful information from OPO reflection. And even after reflection demodulation, we cannot see.
Anyway, we can check more things to confirm.

Participant: Marco, Aritomi, and Yuhang
1. We replaced the OPO transmission PD by PDA10CS. The new PD has a bandwidth of 17MHz so that we can see the oscillation of coherent control error signal(should be 14MHz).
2. Then we brought back BAB and aligned it with OPO scanning and green phase scanning. Here the green phase scanning with 1kHz and 2Vp-p. This parametric amplification effect is not obvious when the modulation magnitude around 1Vp-p. However, we see many fringes within 1ms. This means 2 Vp-p corresponds to the scanning depth is more than 1 period. See attached figure 1.
We also make BAB and p-pol peak overlap when we scanning OPO. We wrote down the p-pol PLL frequency shift. It is 30.5MHz. In this case, green power is 51mW. OPO temperature is 7.03kOm. In principle, this frequency difference will be always like this if we keep green power and OPO temperature. Then we lock p-pol PLL.
3. Then we replace BAB with coherent control beam. Lock PLL with 7MHz. As soon as we did that, we found a phenomenon in the attached picture 2. It seems there is a very clear oscillation at high frequency. After we lock OPO and we look into the detail of this oscillation. We found this oscillation is around 14MHz. See attached picture 3. It seems the coherent control signal we are looking for. However, we found a not understandable error signal after we did the demodulation. Because we found the error signal didn't change oscillation frequency after we changed the green phase scanning frequency. So further investigation needs to be done.
4. We didn't see any useful information from OPO reflection.
We have noticed the nonlinear effect became worse for the second time of CC error signal checking. This proofs that we should consider more about how to apply coherent control.

Participant: Marco and Yuhang
We found a sudden movement of the mirror while we were doing the experiment. We think two mirrors are suspicious, which are used for coherent control laser alignment. Because the alignment of p-pol didn't change, we think the guess above is reasonable.
The suspicious mirrors are marked with black circles in the attachment. Next time, we should try to recover alignment by moving only one mirror. So that we can know which one is causing the problem and replace it.
The fork is properly fixed.